1) Write the mathematical expression for the reaction quotient (Qc) for each of the following reactions:


(a) CH4(g)+ Cl2(g) ⇌ CH3Cl(g) + HCl(g)


(b) N2(g) + O2(g) ⇌ 2NO(g)


(c) 2SO2(g) + O2(g)⇌ 2SO3(g)


(d) BaSO3(s) ⇌ BaO(s) + SO2(g)


(e) P4(g) + 5O2(g) ⇌ P4O10(s)


(f) Br2(g) ⇌ 2Br(g)


(g) CuSO4⋅5H2O(s) ⇌ CuSO4(s) + 5H2O(g)


 

2) Consider the following reaction:


S O 2 C l 2 ( g ) ⇌ S O 2 ( g ) + C l 2 ( g )

At a given point in a reaction,


[SO2Cl2] = 0.12 M, [Cl2] = 0.16 M and [SO2] = 0.050 M. 


Kc = 0.078.


Determine if the following system is at equilibrium. If not, in which direction will the system need to shift to reach equilibrium?


 


3) Nitrogen and oxygen react at high temperatures.


(a) Write the expression for the equilibrium constant (Kc) for the reversible reaction:


N2(g) + O2(g) ⇌ 2NO(g)   ΔH = 181kJ


 


(b) What will happen to the concentrations of N2, O2, and NO at equilibrium if more O2 is added?


(c) What will happen to the concentrations of N2, O2, and NO at equilibrium if N2 is removed?


(d) What will happen to the concentrations of N2, O2, and NO at equilibrium if NO is added?


(e) What will happen to the concentrations of N2, O2, and NO at equilibrium if the pressure on the system is increased by reducing the volume of the reaction vessel?


(f) What will happen to the concentrations of N2, O2, and NO at equilibrium if the temperature of the system is increased?


(g) What will happen to the concentrations of N2, O2, and NO at equilibrium if a catalyst is added?


 


4) Hydrogen is prepared commercially by the reaction of methane and water vapor at elevated temperatures.


CH4(g) + H2O(g) ⇌ 3 H2(g) + CO(g)


What is the equilibrium constant for the reaction if a mixture at equilibrium contains gases with the following concentrations:


CH4, 0.126 M;


H2O, 0.242 M;


CO, 0.126 M;


H2 1.15 M,


at a temperature of 760 °C?


 


5) Analysis of the gases in a sealed reaction vessel containing NH3, N2, and H2 at equilibrium at 400 °C established the concentration of N2 to be 1.2 M and the concentration of H2 to be 0.24 M.


N2(g) + 3 H2(g) ⇌ 2 NH3(g) Kc=0.50 at 400°C


Calculate the equilibrium molar concentration of NH3.


 


6) Calculate the pressures of all species at equilibrium in a mixture of NOCl, NO, and Cl2 produced when a sample of NOCl with an initial pressure of 10.0 atm comes to equilibrium according to this reaction:


 2 NOCl(g) ⇌ 2 NO(g) + Cl2(g)  KP=4.0×10^-4


1
Expert's answer
2021-01-25T04:32:41-0500

1. Qc describes a reaction that is not at the equilibrium.

a)."Qc=\dfrac{[CH_3Cl][HCl]]}{[CH_4][Cl]}"


b.)"Qc=\dfrac{[NO]^2}{[N_2][O_2]}"


c.)"Qc=\dfrac{[SO_3]^2}{[SO_2]^2[O_2]}"


d."Qc=\dfrac{[BaO][SO_2]}{[BaSO_3]}"


e)"Qc=\dfrac{[P_4O_{10}]}{[P_4][O_2]^5}"


f.)"Qc=\dfrac{[Br]^2}{[Br_2]}"


g)"Qc=\dfrac{[CuSO_4][H_2O]^5}{[CuSO_4.5H_2O]}"

2.)

"Qc=\dfrac{[SO_2][Cl_2]}{[SO_2Cl_2]}"

Substituting the given values into the formula;

"Qc=\dfrac{[0.050][0.16]}{[0.12]}"

"Qc=0.067"

Therefore,the equilibrium shifts to the left since the value of "Qc>Kc"


3.

"Kc" for the reversible reaction is

"Kc=\dfrac{[N_2][O_2]}{[NO]^2}"


a. If "O_2" is added the system will shift the equilibrium to the right towards the product side.

b. If "N_2" Is removed the rate of reaction will decrease.

c. If "NO" is added the reverse reaction will be favored.

If catalyst is added it decreases the amount of time taken for the system to reach equilibrium.

4.

"Kc=\dfrac{[H_2]^3[CO]}{[CH_4][H_2O]}"



"Kc=\dfrac{[1.15]^3[0.126]}{[0.126][0.242]}"

"Kc=6.286"